# Non-Gaussian state teleportation with a nonlinear feedforward

### v4

@@ Abstract
Measurement-induced quantum computation with continuous-variable cluster states utilizes teleportation propagating the states through the cluster accompanied by non-Gaussian measurements and feedforward control. We analyze such propagation of a quantum non-Gaussian state with nonlinear squeezing through a small cluster state and show that when a nonlinear feedforward is involved in the teleportation protocol, higher nonlinear squeezing can be transferred. In a probabilistic regime, the improvement can be manifested even with current experimental resources. Better processing of non-Gaussian states can bring us closer to the necessary interplay between cluster states and non-Gaussianity required by quantum computing.
@@ 1 Introduction
Entangled Gaussian cluster states offer a rich multimode structure. Accompanied by single photon measurements, their simulation complexity exceeds the abilities of classical computation [GBS, Chabaud2023]. Their capacity for massive multi-mode multiplexing [OBrien2007, Yokohama2013, Larsen, Menicucci] is the main advantage of measurement-induced quantum computing with bosonic modes of traveling light. In this protocol, measurements of the individual modes enable the quantum information to propagate along the cluster and evolve in the desired manner [ClusterCV, Menicucci, Walshe2020]. The choice of measurement is what determines whether the state is simply transmitted by quantum teleportation [Ban2002, Furusawa1998, Pirandola2015, Walshe2020, Takeoka] or transformed [Sefi2019, Filip2005, Miyata2016, Hillmann2022]. To reach the full advantages of the quantum computation, the Gaussian entangled states need to be further supported by deterministic non-Gaussian operations [Lloyd1999, Gottesman2001, Mari2012, Chabaud2023]. These challenging operations can be implemented in a measurement-induced fashion, [Miyata2016, Konno2021b] leaving it with two distinct tasks: experimental generation of the non-Gaussian states, and successful propagation of the non-Gaussian states through a Gaussian circuit.
@@
In quantum optics with traveling light, non-Gaussian quantum states have been generated in a wide range of experimental configurations [Yukawa2013b, URen2004, Ogawa2016, Zavatta2004, Ourjoumtsev2006, LachmanG19]. Today, photon-number-resolving detectors, together with Gaussian operations such as parametric down-conversion and coherent classical displacements, are feasibly used to probabilistically generate non-Gaussian states with varied nonclassical and non-Gaussian properties [Konno2021, Takeoka2011, Asavanant17, Melalkia22, Tiedau, Konno2024]. One of these important properties is the nonlinear squeezing that can be used as a resource for deterministic preparation of cubic nonlinearity [Miyata2016, sakaguchi] and is a specific example of a purely continuous non-Gaussian resource [Kala22, Konno2021, Brauer2021]. The basic principle of the state preparation, closely aligned to the Gaussian boson sampling problem [GBS], is still actively investigated to prepare more and more complex quantum states [Takase2023]. The successful generation of non-Gaussian states led to their active use in tests of their propagation through Gaussian circuits, resulting in demonstration of key protocols, such as teleportation [Ide, Mista, Fuwa, Takeda, Benichi, Lee, Zhao2023], decoherence prevention [Jeannic2018], or nonlinear measurement [sakaguchi]. One thing shared by all of these experimental tests is that the non-Gaussian features of the employed states are deteriorated in the process, diminishing their power. Therefore, novel ways of processing that help preserve the non-Gaussianity will be needed while moving forward.
In this paper we are going to show how the canonical teleportation protocol along a single node of an entangled cluster can be enhanced by a classical non-linear feedforward to better transfer a specific class of non-Gaussian quantum states. We focus on quantum states with cubic nonlinear squeezing which is a resource for measurement-based quantum computation with continuous variables. To this end we are going to consider the problem of teleporting these states, substantially different than Fock states, qubit states and cat states, and compare the canonical teleportation to the teleportation with enhanced feedforward operation. In the field of measurement-induced operations, the teleportation of a state with nonlinear squeezing enriched by nonlinear feedforward can be seen as an approximation of the cubic nonlinear phase measurement [sakaguchi]. The teleported state serves as an ancillary state and together with the homodyne measurements and feedforward they construct the approximated measurement which applies to the first mode of the two-mode Gaussian state. Therefore, our analyses will also test the projection powers of the approximate measurement as a state preparation. We are going to compare this approximation to the ultimate ideal theoretical bound provided by the ideal cubic measurement.
@@ 2 Teleportation, linear and nonlinear
2.1 Deterministic regime
The teleportation protocol [Furusawa1998, Pirandola2015] is the basic building block of the cluster state computing. The usual involved measurement is the homodyne detection with a linear feedforward that provides linear dependence of operations performed on the outcoming state on measurements results. It is possible to consider also generalized measurement, that enables not only transfer of quantum states but also their transformation. Such a measurement can include adjustable nonlinear feedforward that utilizes the measurement results in some more complex function.
Figure 1: a) Regular teleportation scheme transferring the quantum state through the cluster state. b) Optical scheme equivalent to parameterization of the cluster state considered for the analyzed schemes. c) Optical scheme showing the nonlinear cubic phase measurement applied to one part of the cluster state. The measurement is decomposed into an ideal cubic nonlinearity followed by homodyne measurement. The measurement result is used in a feedforwarded displacement on the second mode. d) Nonlinear teleportation with nonlinear feedforward and ancillary state in the mode Q.
A single step of the transfer, the canonical teleportation protocol itself, is schematically depicted in Fig. 1a. Within the protocol, the initial state in mode $Q$ interacts on a balanced beam splitter with mode $A$ of a two-mode entangled state and the two resulting modes are measured by a pair of homodyne detectors. The measured values are then used for the feedforward displacement operation on the remaining mode $B$ of the entangled state, leaving its quadrature operators transformed as
The expressions used for individual transformations generated by beam splitter and other optical elements can be found in Appendix A. Quadratures $x$ and $p$ of the optical field obey $[x,p]=i$. The statistics of the teleported state is accompanied by noise terms $N_{x}$ and $N_{p}$, that depend on the properties of the two-mode entangled state, illustrated in Fig. 1b, and vanish for the ideal EPR state with maximal correlation and infinite energy. Despite the protocol yielding perfect results in the limit of infinite squeezing, the real performance will always be noisy, with an imperfect outcome.
Our goal is teleportation of a quantum non-Gaussian state. Specifically the cubic phase state
@@ 2.1 Deterministic regime
where
$C(\chi)=\exp(-i\chi\frac{x^{3}}{3}),$ (4)
@@
$O=p+zx^{2},$ (5)
that projects on its eigenstates. Such states have infinite energy, similar to position eigenstates, which are projected on by homodyne measurement [Gottesman2001, Ghose2007]. The measurement can be decomposed into the cubic nonlinearity and p-homodyne measurement. These operations can be sequentially applied to one part of the two-mode cluster state as is shown in Fig. 1c. It is necessary to utilize the measurement result $m$ in a Gaussian feedforward in order to obtain a non-Gaussian state. If the measurement outcome is disregarded, the outcoming state is quantum Gaussian, i.e. it is a mixture of Gaussian states. In Heisenberg picture, the overall transformation yields
@@
Here the output is displaced in $p$ over the measurement result with gain $g_{p}$. Note the appearance of nonlinear terms that originates from the cubic nonlinearity, which action is in more detail described in Appendix A.
@@
Whereas the noise term $N_{x}$ remains the same as in the case of regular teleportation, new nonlinear terms appear in transformation of the $p$ quadrature
@@
Therefore, the $p_{Q}$ quadrature of the ancillary mode Q that initially carries the nonlinear squeezing is nonlinearly transformed and accompanied by non-Gaussian noise terms.
@@
with $d_{F}$ and $g_{F}$ being electronic gains, and $t$ and $r$ being the transmissivity and reflectivity of the beam splitter before the homodyne measurements. By setting $\chi=0$, the equation (9) describes the regular teleportation with generalized gains and beam splitter transmissivity.
@@ 2.2 Probabilistic regime
Schemes described in the previous subsection work in the deterministic regime, accepting all measurement results and processing them in the feedforward. Different measurement results of the homodyne measurements $m_{x}$ and $m_{p}$ can generally lead to a different quantum states $\hat{\rho}(m_{x},m_{p})$
@@
Here $\rho_{\textrm{ABQ}}$ is the density matrix of the cluster state and ancillary mode and $\Pi(m_{x},m_{p})$ a POVM element describing the measurement. For the ideal nonlinear measurement shown in Fig. 1c, it is the POVM in (3). For the regular and nonlinear teleportation it yields
@@
If the integration domain $r$ is the whole phase space, this is the state, whose nonlinear squeezing is computed in the deterministic regime. Generally, the states $\rho(m_{x},m_{p})$ are different and thus, it can be advantageous to post-select some of them (i.e. choose $r$ as an appropriate subspace of the phase space) in order to increase the nonlinear squeezing at the output. This is, however, accompanied by decreased probability of success.
@@
The analysis is performed in a following way. The density matrices of quantum states are computed for measurement results on a grid in phase space, such that trace of their mixture approaches very closely one. Thus, the grid is sufficiently large to capture all states appearing with numerically significant probability. Then, the states are ordered according to their nonlinear squeezing and aggregated up to some probability $P$. For this mixture of post-selected states, the nonlinear squeezing is evaluated.
@@
As a preparation of a general two mode Gaussian state is quite complex, we conclude with analysis of a linear unity gain teleportation through a two mode squeezed vacuum compared with the same setup enriched by the nonlinear feedforward. In fact, the nonlinear feedforward can be more complex, also adjusting the outcoming states $\rho(m_{x},m_{p})$ by Gaussian operations in such a way, that the nonlinear squeezing is maximized for each of the state. This is equivallent to computation of averaged nonlinear squeezing, where not the conditioned states are aggregated, but their nonlinear squeezing
$\expectationvalue{\xi(z)}=\int_{r}P(m_{x},m_{p})\xi(z)_{\rho(m_{x},m_{p})}dm_{x}dm_{p}.$ (16)
@@
In such a way we present comparison of straightforward scheme of unity gain teleportation through a two mode squeezed vacuum with and without the nonlinear feedforward.
2.3 Nonlinear squeezing
Both the regular and nonlinear teleportation schemes consists of only Gaussian optical elements, therefore, with Gaussian states at the input, the circuit yields classical non-Gaussian mixture of Gaussian states. Any non-Gaussianity of the output state therefore has to originate in the non-Gaussianity of the input state. Our goal here is to look at the teleported state and compute how much of its non-Gaussian properties survived the teleportation
The original role of non-Gaussian ancillary state is to suppress a noise term that appears at the outcome of the measurement-induced cubic phase gate [Miyata2016, Konno2021]. This term has the same form as the operator (5) that is measured by the nonlinear measurement; the real parameter $z$ is called cubicity. Inspired by insufficiency of Gaussian states for the quantum computing advantage with CV, the nonlinear squeezing is defined as a ratio of variance of the operator $O$ compared to the minimal variance that can be achieved in a Gaussian state [Kala22, Brauer2021, Konno2021]
As a consequence of this definition a state with nonlinear squeezing is certainly non-Gaussian. Our goal, from now on, will be to analyze how well a quantum state with this property can be transferred through a two-mode cluster state. We will specifically look for the maximal nonlinear squeezing attainable at the output of the teleportation protocol with a specific cubic nonlinear state at its input. Subsequently, we will compare this results to the teleportation protocol enhanced by a nonlinear feedforward. Thus, the figure of merit will be the value of the nonlinear squeezing at the output.
The maximal nonlinear squeezing is reached by the unphysical infinite energy cubic state (18). Such a state is well approximated by a cubically transformed squeezed vacuum
[…]
This finitely nonlinearly squeezed state has finite energy and minimizes uncertainty relations. However, it is an idealized approximation in a sense that its preparation still includes experimentally unfeasible cubic nonlinearity.
Experimentally feasible approximations of the finitely nonlinearly squeezed states can be prepared on a finite Fock subspace [Yukawa2013b, Huang]. The approximations can be written as
[…]
where the coefficients can be found in an efficient way by solving an eigenvalue problem and two variable optimization [Miyata2016]. Note that Fock states and states with nonlinear squeezing have different symmetries. Unlike the nonlinearly squeezed states that are symmetric in $x$ only, Wigner functions of Fock states are symmetric both in $x$ and $p$. This difference hints that an approximation of the cubic state must be a superposition of Fock states. The experimentally feasible quantum state exhibiting nonlinear squeezing and simultaneously the lowest order approximation is a superposition of the first two Fock states
[…]
where $u$ is real and bounded by $0<u<1$ [Konno2021]. The used parameterization ensures symmetry under transformation $x\rightarrow-x$ which holds also for (5). One step further leads to ancillary state approximation
[…]
where the $c_{i}$ are real coefficients. The states can be found as eigenstates of the operator $O$ restricted to an $N$ dimensional subspace of the Fock space. They provide the best nonlinear squeezing available on such a subspace [Miyata2016].
To show the positive effect of nonlinear feedforward, let us compare the results obtained by the regular and nonlinear teleportation protocols. At the same time, we compare the schemes equipped with identical resources of non-Gaussianity [Takagi] and Gaussian squeezing [Braunstein2005b]. The approximations can be classified by the faithful hierarchy of $N$-photon non-Gaussianity [LachmanG19] or stellar rank [chabaud2020]. A superposition (19) with a fixed highest Fock state $N$ has $N$-photon non-Gaussianity. When teleporting a state with some given $N$-photon non-Gaussianity and transferring a state from the same level of the hierarchy by nonlinear teleportation, both schemes posses the same $N$-photon non-Gaussian state and any improvement in performance thus originates from the properties of the scheme. Another resource present in the schemes is Gaussian squeezing [Braunstein2005b]. Throughout our analysis we limit the available Gaussian squeezing while comparing the performances. At first in preparation of the finitely squeezed state (18) and as well of the cluster state. We consider a general two-mode Gaussian cluster state, whose parameters are optimized for each situation and each scheme. The optimization is done with help of Python library scipy optimize and repeated with randomly chosen initial conditions. Details can be found in Appendix C. To remain close to the experimental reality, we consider an upperbound on Gaussian squeezing available. The optimum is then sought in a range given by these bounds, e.g. $\pm$9 dB. Overall, by limiting the resources we can perform a genuine comparison of schemes equipped with the same level of $N$-photon non-Gaussianity and limit on maximal Gaussian squeezing.
The parameterization of the two-mode Gaussian state via the covariance matrix can be found in Appendix B and is equivalent to two thermal states transformed by single mode Gaussian operations and a beam splitter as shown in Fig. 1b. The thermal noise is taken as an input parameter and by its change, we can analyze the resilience of discussed effects. It is described by parameter $n$ representing the multiples of vacuum fluctuations added to an initially pure vacuum state. Thus, for a pure state the parameter yields $n=0$ , while $n>0$ represents some amount of thermal noise [Mattia2020].
When processing a quantum state or seeking nonlinear squeezing after a nonlinear measurement on an entangled mode, the nonlinear squeezing of the output state does not need to be minimal for the initial cubicity. Thus we define the native cubicity $z_{n}$ of a given state $\rho$ as the cubicity for which
$\xi(z_{n})_{\rho}=\min_{z}\xi(z)_{\rho}.$ (22)

### v5

@@ Abstract
Measurement-induced quantum computation with continuous-variable cluster states utilizes teleportation to transmit and alter quantum states via measurement-and-feedforward control. One of the key challenges of this approach is the deterioration of quantum states caused by the noise added due to imperfect entanglement of the cluster. We analyze the propagation of a quantum non-Gaussian state with nonlinear squeezing through a small cluster state. We show that a nonlinear feedforward in the deterministic teleportation protocol reduces the added noise and improves the nonlinear squeezing transferred. In a probabilistic regime, the improvement can be manifested even with current experimental resources. Better processing of non-Gaussian states can bring us closer to the necessary interplay between cluster states and non-Gaussianity required by quantum computing.
@@ 1 Introduction
The capacity of Gaussian cluster states for massive multi-mode multiplexing [36, 55, 26, 33] is the main advantage of measurement-induced quantum computing with bosonic modes of traveling light. Accompanied by non-Gaussian measurements, their simulation complexity exceeds the abilities of classical computation [16, 9]. The main underlying working principle of the processing with the Gaussian cluster states is quantum teleportation. Measurements of the individual modes enable manipulating the quantum information as it propagates along the cluster [32, 33, 53]. The choice of measurement determines whether the state is simply transmitted by quantum teleportation [2, 12, 41, 53, 48] or transformed in a required manner [44, 11, 35, 17]. To reach the full advantage of the quantum computation, it is necessary to employ deterministic non-Gaussian operations [29, 15, 30, 9, 38] along with the Gaussian entangled cluster state. Experimental challenges motivated measurement-induced implementation of those operations, [35, 22] leaving two distinct tasks: experimental generation of the non-Gaussian states and their successful propagation through a Gaussian circuit.
@@
In quantum optics with traveling light, non-Gaussian quantum states have been generated in a wide range of experimental configurations [56, 51, 37, 57, 40, 25]. Today, photon-number-resolving detectors, together with Gaussian operations such as parametric down-conversion and coherent classical displacements, are used to probabilistically generate non-Gaussian states with varied nonclassical and non-Gaussian properties [24, 49, 1, 31, 50, 23]. The basic principle of the state preparation, closely aligned to the Gaussian boson sampling problem [16], is still actively investigated to prepare more complex quantum states [46].
The successful generation of non-Gaussian states led to their active use in tests of their propagation through Gaussian circuits, resulting in the demonstration of key protocols. One of the most fundamental is quantum teleportation [19, 34, 13, 47, 4, 27, 58], which is the cornerstone of measurement-based quantum computing [33]. The key challenge it is facing is the deterioration of quantum states due to the noise caused by imperfect quantum resources and as well experimental imperfections. This deterioration can be alleviated by adaptation - either of the protocol [10], or of the quantum states [20, 21, 42]. The crucial decision that needs to be made is which of the features of a quantum state should be protected, as it is too challenging to protect all. Fidelity and Wigner function negativity enable to show average quality and necessary non-Gaussianity of a state processed by the quantum protocol, however, they are not sufficiently operational. Here we use the more practical notion of nonlinear squeezing. It appears as a perfect choice due to its operational definition, clear and straightforward interpretation in continuous variable quantum information protocols, and relevance for current experiments. As such, it enables us to proceed towards the concepts defined with ideal resources [22, 3].
The nonlinear squeezing is defined as a noise suppression in a nonlinear function of quadratures that falls below a Gaussian threshold. The nonlinear squeezing is a resource for the deterministic realization of cubic phase gate [35], which appears in the continuous variable quantum information in two contexts, as a part of the non-Clifford operation on GKP qubits [15] and as an operation unlocking a universal control of a continuous mode [29]. When a measurement-induced cubic operation is implemented via cubic nonlinear phase measurement [43], the nonlinear squeezing is a single property required from an auxiliary state. It directly determines the amount of unwanted noise added to the target result. Simultaneously, it is a specific example of a purely continuous-variable sign of non-Gaussianity [35, 21, 5], relevant to current experimental techniques [24] and different to e.g. Fock or Schrödinger cat states [39, 50].
In this paper, we study the measurement-induced propagation of the cubic nonlinear squeezing along a single node of a cluster state. We consider the nonlinear squeezing as both the resource of non-Gaussianity and the figure of merit characterizing the processed state. We evaluate the performance of the canonical teleportation protocol optimized for nonlinear squeezing transfer and show its limits. Importantly, we also show that it is possible to go beyond these limits when the teleportation is enhanced by using the nonlinear feedforward [35, 43]. Finally, we show the ultimate limits on teleportation of cubic squeezing provided by remote state preparation by ideal cubic measurement. All these results provide a vital insight, necessary for further extensions of such measurement-induced protocols with large cluster states.
@@ 2 Teleportation, linear and nonlinear
2.1 Nonlinear squeezing
The original role of the non-Gaussian auxiliary state with the cubic nonlinear squeezing is to suppress a noise term that appears at the outcome of the measurement-induced cubic phase gate [35, 24]
$O=p+zx^{2},$ (1)
the real parameter $z$ is called cubicity. Inspired by the insufficiency of Gaussian states for the quantum computing advantage with CV, the nonlinear squeezing is defined as a ratio of the variance of the operator $O$ compared to the minimal variance that can be achieved in a Gaussian state [21, 5, 24]
As a consequence of this definition, a state with nonlinear squeezing ($\xi(z)<1$) is certainly non-Gaussian. The nonlinear squeezing can also be expressed in dB as
$\xi(z)[\textrm{dB}]=10\log_{10}(\xi(z)).$ (3)
Our goal, from now on, will be to analyze the quality of the nonlinear squeezing transferred through the cluster state. Simultaneously, the nonlinear squeezing injected into the cluster state serves as the only source of non-Gaussianity here. Its evaluation at the output is a figure of merit of the procedure quality.
The maximal nonlinear squeezing is reached by the unphysical infinite energy cubic state
@@ 2.1 Nonlinear squeezing
$C(\chi)=\exp(-i\chi\frac{x^{3}}{3}).$ (5)
Such a state is well approximated by a cubically transformed squeezed vacuum
@@
where the squeezed vacuum approximates the quadrature eigenstate. This finitely nonlinearly squeezed state has finite energy and minimizes uncertainty relations. However, it is an idealized approximation in the sense that its preparation still includes experimentally unfeasible cubic nonlinearity.
Experimentally feasible approximations of the finitely nonlinearly squeezed states can be prepared on a finite Fock subspace [56, 18]. The approximations can be written as
@@
where the coefficients can be found efficiently by solving an eigenvalue problem and two variable optimization [35]. Note that Fock states and states with nonlinear squeezing have different symmetries. Unlike the nonlinearly squeezed states that are symmetric in $x$ only, the Wigner functions of Fock states are symmetric in $x$ and $p$. This difference hints that an approximation of the cubic state must be a superposition of Fock states. The experimentally feasible quantum state exhibiting nonlinear squeezing and simultaneously the lowest order approximation is a superposition of the first two Fock states
[…]
where $u$ is real and bounded by $0<u<1$ [24]. The used parameterization ensures symmetry under transformation $x\rightarrow-x$ which holds also for (1). One step further leads to auxiliary state approximation
[…]
where the $c_{i}$ are real coefficients. The states can be found as eigenstates of the operator $O$ restricted to an $N$ dimensional subspace of the Fock space. They provide the best nonlinear squeezing available on such a subspace [35]. In the further analysis, we employ those states as the resource for non-Gaussian measurements and operations. When processing a quantum state or seeking nonlinear squeezing after a nonlinear measurement on an entangled mode, the nonlinear squeezing of the output state does not need to be minimal for the initial cubicity. Thus we define the native cubicity $z_{n}$ of a given state $\rho$ as the cubicity for which
$\xi(z_{n})_{\rho}=\min_{z}\xi(z)_{\rho}.$ (10)
2.2 Deterministic regime
The teleportation protocol [12, 41] is the basic building block of cluster state computing. Usually, a homodyne measurement is used together with a linear feedforward. The outcoming state is processed via a Gaussian displacement operation that linearly depends on the homodyne measurement results. Following the idea of cluster state computing, it is possible to consider also generalized measurements, that enable not only the transfer of quantum states but also their transformation [11, 33]. Such measurements can include an adjustable nonlinear feedforward that utilizes the measurement results in some more complex function.
Figure 1: a) Optical scheme illustrating the considered cluster state generation. b) Optical scheme showing the nonlinear cubic phase measurement applied to one part of the cluster state to project it onto the cubic state. The measurement is decomposed into an ideal cubic nonlinearity followed by homodyne measurement. The measurement result is used in a feedforwarded displacement on the second mode. c) Regular teleportation scheme with linear feedforward transferring the quantum state through the optical cluster state. d) Nonlinear teleportation with nonlinear feedforward and auxiliary state in the mode Q.
A single step of the transfer, the canonical teleportation protocol itself, is schematically depicted in Fig. 1c. The initial state in mode $Q$ interacts on a balanced beam splitter with mode $A$ of a two-mode entangled state. The modes $Q$ and $A$ are measured by a pair of homodyne detectors. The measured values are then used for the feedforward displacement operation on the remaining mode $B$ of the entangled state. Altogether, the canonical teleportation protocol transforms the initial quadrature operators as
The expressions for individual transformations generated by beam splitter and other optical elements can be found in Appendix A. Quadratures $x$ and $p$ of the optical field obey $[x,p]=i$. The statistics of the teleported state are accompanied by noise terms $N_{x}$ and $N_{p}$, which depend on the properties of the two-mode entangled state, illustrated in Fig. 1a, and vanish for the ideal EPR state with maximal correlation and infinite energy. Despite the protocol yielding perfect results in the limit of infinite squeezing, the real performance will always be noisy, with an imperfect outcome.
Our goal is to obtain the nonlinear squeezing at the mode $B$ of the cluster state. To this end, we replace the canonical teleportation protocol with the nonlinear phase cubic measurement [43] and, moreover, its ideal theoretical limit of projection onto the cubic state. The POVM element of the projection can be expressed as
[…]
@@ 2.2 Deterministic regime
Note the appearance of nonlinear terms that originate from the cubic nonlinearity, which action is in more detail described in Appendix A.
@@
Whereas the noise term $N_{x}$ remains the same as in the case of regular teleportation, new nonlinear terms appear in the transformation of the $p$ quadrature
@@
The $p_{Q}$ quadrature of the auxiliary mode Q that initially carries the nonlinear squeezing is nonlinearly transformed and accompanied by non-Gaussian noise terms. The transformation already includes the nonlinear feedforward that influences the expression via the parameter $\chi$.
@@
with $d_{F}$ and $g_{F}$ being electronic gains, and $t$ and $r$ are the transmissivity and reflectivity of the beam splitter before the homodyne measurements. By setting $\chi=0$, the equation (16) describes the regular teleportation with generalized gains and beam splitter transmissivity.
To show the positive effect of nonlinear feedforward, let us compare the results obtained by the canonical and nonlinear teleportation protocols, Fig. 1c and d. At the same time, we compare the schemes equipped with identical resources of non-Gaussianity [45] and Gaussian squeezing [6]. The approximations can be classified by the faithful hierarchy of $N$-photon non-Gaussianity [25] or stellar rank [8]. A superposition (7) with a fixed highest Fock state $N$ has $N$-photon non-Gaussianity. When teleporting a state with some given $N$-photon non-Gaussianity and transferring a state from the same level of the hierarchy by nonlinear teleportation, both schemes possess the same $N$-photon non-Gaussian state , and any improvement in performance thus originates from the properties of the scheme. Another resource present in the schemes is Gaussian squeezing [6]. Throughout our analysis, we limit the available Gaussian squeezing while comparing the performances. At first in preparation of the finitely squeezed state (6) and of the cluster state as well. We consider a general two-mode Gaussian cluster state, whose parameters are optimized for each situation and each scheme. Specificaly, we optimize independently $r_{1},r_{2},\phi,\phi_{1},\phi_{2},\alpha_{1}$ and $\alpha_{2}$ in the preparation of the cluster state as illustrated in Fig. 1a. Further, we optimize all gains and transitivities of beam splitters in schemes Fig.1b,c and d. Eventually, we optimize all free parameters of the teleported/ auxiliary states $\chi,r,u,c_{0},c_{1}$ and $c_{2}$ in (6),(8) and (9) The optimization is done with help of Python library scipy optimize and repeated with randomly chosen initial conditions. Details can be found in Appendix C. To remain close to the experimental reality, we consider an upper bound on Gaussian squeezing available. The optimum is then sought in a range given by these bounds, e.g. $\pm$9 dB. Overall, by limiting the resources we can perform a genuine comparison of schemes equipped with the same level of $N$-photon non-Gaussianity and limit on maximal Gaussian squeezing.
The parameterization of the two-mode Gaussian state via the covariance matrix can be found in Appendix B and is equivalent to two thermal states transformed by single mode Gaussian operations and a beam splitter as shown in Fig. 1a. The thermal noise is taken as an input parameter and by its change, we can analyze the resilience of discussed effects. It is described by a parameter $n$ representing the multiples of vacuum fluctuations added to an initially pure vacuum state. Thus, for a pure state the parameter yields $n=0$ , while $n>0$ represents some amount of thermal noise [52].
@@ 2.3 Probabilistic regime
Schemes described in the previous subsection work in the deterministic regime, accepting all measurement results and processing them in the feedforward. Different measurement results of the homodyne measurements $m_{x}$ and $m_{p}$ can generally lead to different quantum states $\hat{\rho}(m_{x},m_{p})$
@@
Here $\rho_{\textrm{ABQ}}$ is the density matrix of the cluster state and the teleported mode and $\Pi(m_{x},m_{p})$ a POVM element describing the measurement. For the ideal nonlinear measurement shown in Fig. 1b, it is the POVM in (12). For the regular and nonlinear teleportation, Fig. 1c and d, it yields
@@
The deterministic scheme is obtained for $\sigma$ being the whole phase space. Generally, the conditionally prepared states $\rho(m_{x},m_{p})$ are different, and thus, it is advantageous to post-select some of them (i.e., choose $\sigma$ as an appropriate subspace of the phase space) in order to increase the nonlinear squeezing at the output. This is, however, accompanied by a reduction in the probability of success.
@@
The analysis is performed in the following way. The density matrices of quantum states $\rho(m_{x},m_{p})$ are computed for measurement results on a grid of $m_{x}^{i}$ and $m_{p}^{i}$ with steps $\Delta m_{x}$ and $\Delta m_{p}$, such that the trace of their mixture approaches very closely one. Thus, the grid is sufficiently large to capture all states appearing with numerically significant probability. Then, the states $\rho(m_{x}^{i},m_{p}^{i})$ are ordered according to their nonlinear squeezing and aggregated up to some probability $P$. For this mixture of post-selected states, the nonlinear squeezing is evaluated. The corresponding integration domain is $\sigma=\cup_{i}[(m_{x}^{i},m_{x}^{i}+\Delta m_{x})\times(m_{p}^{i},m_{p}^{i}+\Delta m_{p})]$.
@@
To conclude with a straightforward example, we consider unity gain linear and nonlinear teleportation via two-mode squeezed vacuum (5). In fact, the nonlinear feedforward can be more complex, also adjusting the outcoming states $\rho(m_{x},m_{p})$ by Gaussian operations in such a way, that the nonlinear squeezing is maximized for each of the states. This is equivalent to the computation of averaged nonlinear squeezing, where not the conditioned states are aggregated, but their nonlinear squeezing
$\expectationvalue{\xi(z)}=\int_{\sigma}P(m_{x},m_{p})\xi(z)_{\rho(m_{x},m_{p})}dm_{x}dm_{p}.$ (23)
@@
In such a way, we present a comparison of the straightforward scheme of unity gain teleportation through a two mode squeezed vacuum with and without the nonlinear feedforward.